OpenStudy (anonymous):

can someone please help me with this question. What is the solution to the rational equation? x/5 + 4/5 = 7

OpenStudy (zepp):

If you pass the 4/5 to the left side, you would get x/5 = 7 - 4/5

OpenStudy (zepp):

Then you would get something like x/5 = 31/5 x = (31/5)*5

OpenStudy (anonymous):

5(x/5 +4/5)=7(5) x+4=35 x=35-4 x=31

OpenStudy (anonymous):

thank you guys :]